OpenStudy (anonymous):

-3/5 ÷ -6/7 = Answer 18/35 7/10 10/7 35/18

OpenStudy (anonymous):

@matricked

OpenStudy (anonymous):

you can multiply by the reciprocal (-3/5)(-7/6) = 21/30 = 7/10

OpenStudy (anonymous):

7/10